# Learning Object Compliance via Young’s Modulus from Single Grasps using Camera-Based Tactile Sensors

### v2

@@ 4 Physical Dataset
A novel physical dataset of 285 objects was gathered to train and validate our method. For comparison, our dataset is nearly 4 times larger than the YCB dataset, which includes only 77 objects [51]. Chosen objects were intended to be among things a robotic manipulator may commonly interact with in general operation. Additionally, each object was required to have approximately uniform material composition to ensure that an accurate Young’s modulus could be defined for the object and used to label data for supervised training. A subset of dataset objects is displayed in Fig. 4.
@@
Figure 4: Physical dataset. (a) Selected objects from our physical dataset are displayed. These are common objects with a wide variety of shapes and materials. (b) Some custom objects are molded out of silicone to replicate the shape of non-rubber objects in the dataset.
@@
To evaluate how well estimation algorithms could generalize, diversity of shape and material was prioritized in acquiring objects. Thus, we mold 20 custom objects out of silicone to replicate the exact shape of other objects in the dataset with vastly different material composition. Molded objects are displayed in Fig. 4b. Training on data collected across object replicas of different materials will help prevent overfitting to shapes and textures. Moreover, there is incentive to bias towards rubber objects because it is clear from Eq. 5 that it will be easiest to discriminate between Young’s moduli nearest to our sensor’s. The dataset’s composition is further detailed in Appendix E.
As objects were gathered, they were labeled with their Young’s moduli. For harder objects, like those made of plastics and metals, this was done by referencing an engineering database of known material properties [52]. For softer objects, like those made of rubber, Shore hardness was measured and converted to Young’s modulus using Gent’s model of hardness and other established methods [41, 53]. These models provide a reliable measure of object compliance on a logarithmic scale. The dataset includes materials ranging from foam to steel, with Young’s moduli from 5.0 kPa to 250 GPa. For reference, our sensor’s Young’s modulus is 0.275 MPa.
5 Experiments
5.1 Hardware Setup
Figure 5: Hardware setup for data collection. GelSight Wedge sensors are attached to Franka Panda gripper with embedded force gauge.
GelSight Wedge sensors [17] were rigidly attached to a standard parallel robotic gripper with 3D-printed plastic fingers, as depicted in Fig. 5. Although we choose GelSight sensors for our experiment, other flat, camera-based tactile sensors capable of tracking depth in high-resolution would be suitable. One sensor is inscribed with markers. As explained in Appendix B, markers help enhance the dynamic information in a tactile image by tracking surface displacement. Data from only one sensor is required for estimation, but we collect data with both sensors simultaneously to later assess the necessity of markers in estimating Young’s modulus. Additionally, we assume symmetry between fingers in our analytical models. Normal force is measured through a gauge embedded in the left finger. Gripper position is directly measured from control software.
5.2 Data Collection & Training
Figure 6: Raw force and depth data. Data is collected across grasps for spheres of each material type and fit to $d\sim F^{\frac{2}{3}}$, per our Hertzian model.
Data Collection: Automated parallel grasps are executed on objects in fixed position. For each object in our dataset, a set number of grasps are recorded. Objects are re-oriented between each grasp to diversify potential antipodal grasp locations. During grasping, tactile images from GelSight sensors, forces, and gripper widths, are recorded at 30 Hz. Measurements are synchronized by shifting for latency between cameras. Grasps are executed at constant velocity up to force of 30 N. Raw data collected from grasps of spherical objects is plotted in Fig. 6. More details about processing data collected from grasps is provided in Appendix C.
Tactile and depth images at grasping force of 30 N for a set of example objects are depicted in Fig. 7. We observe that textures are more finely captured for rigid object’s with higher modulus. As expected by Eq. 4, peak depth is higher for objects of the same shape with higher modulus.
Data Preprocessing: Data recorded from grasping objects is clipped to include only the loading sequence from first contact to peak force. Misaligned grasps where loading is non-monotonic are disregarded. From here, a set number of frames $N_{\text{frames}}=3$ are sampled equidistantly in time. Multiple augmentations of samples are taken from each grasp. As depicted in Fig. 2, tactile images are sent as inputs to our multi-tower architecture alongside analytical estimates and force and width measurements. Images are randomly transformed with horizontal and vertical flipping.
@@ 5.2 Data Collection & Training
Figure 7: Example tactile data. Tactile images and depth reconstruction from a GelSight Wedge sensor without markers over a set of example objects. Shown frames are all from peak pressing force of 30 N. Tactile images are size 35mm by 25mm.
@@
Training: We train our model for 80 iterations over 4,000 grasp inputs. In total, these grasps include more than 12,000 tactile images. Output labels are normalized on a log10 scale given the range of Young’s moduli in our dataset. Final sigmoid activation function is used with an MSE loss function.
6 Experimental Results
Figure 8: Young’s modulus predictions for rubber objects only. Results for model trained on only rubber objects over 80 randomly sampled grasps of unseen objects. Predictions in the gray region are considered sufficiently accurate.
@@ 6 Experimental Results
Initially, we train an instance of our model using only rubber objects from our dataset. From Eq. 5, we expect it will be easier to discern fine differences in compliance between rubber objects, as their Young’s modulus is closest to that of our sensor. Estimations across unseen objects for this model are plotted in Fig. 8. The model demonstrates 100% prediction accuracy within an order of magnitude of ground truth Young’s modulus, with an average log10 error of 0.36. We utilize this reduced training configuration to compare our architecture against previously developed Shore hardness models [14]. As detailed in Appendix D, our hybrid system exhibits reduced error over this limited domain.
Subsequently, we train our neural network over the entire collected dataset and compare against analytical baselines. Predictions across unseen objects for each method are plotted in Fig. 9. A traditional Hooke’s Law method without tactile sensing fails to distinguish between material types. Simple elastic and Hertzian analytical models exhibit stronger performance, but remain less than 30% accurate. Our hybrid learned model showcases dramatic improvements over all methods.
In Table 1, performance is compared between different estimation methods and inputs. Performance metrics are computed for grasps of both seen and unseen objects. We evaluate models on a logarithmic scale given the range of Young’s moduli in our dataset. We compute error as the average difference between prediction and ground truth on a log10 scale. Predictions are considered accurate within the same order of magnitude of ground truth labels. This tolerance is deemed sufficient for downstream tasks of produce sorting based on the change in compliance of selected fruits and vegetables during ripening [54, 55, 56, 57].
@@
Figure 9: Young’s modulus predictions by material. Results for learned and analytical estimation methods are plotted, using predictions from 80 randomly sampled grasps per material category. Predictions in the gray region are considered sufficiently accurate.
From the ablation study, we observe that our hybrid model drastically out-performs purely analytical and learning-based methods. RGB tactile images with markers are shown to be most favorable for this estimation task. There is not significant difference in results between models which consider force $F$ and width $w$ measurements. This may be due to the fact that all objects were grasped with identical peak force. Moreover, force is implicitly apparent through tactile difference images [12]. Importantly, analytical estimates $\hat{E}_{\text{elastic}},\hat{E}_{\text{hertz}}$ are shown to improve prediction accuracy by nearly 10%. This indicates that a hybrid approach to compliance estimation is favorable over purely learning-based methods, which have been previously explored [14, 15, 1, 2, 3, 4]. Prediction accuracy across shape and material categories for each method is investigated in Appendix E. We observe that our model performs independent of shape and can generalize well across materials.
@@
Table 1: Estimation Architecture Ablation Study
@@
Limitations: Our estimation system struggled to express precise differences in Young’s modulus between harder objects that stray further from the modulus of our sensor (i.e. plastic, wood, and metal objects). These results are expected given that harder objects are rigid relative to our soft sensor. The resolution of our tactile sensor is not high enough to detect minute differences in deformation between this range of compliances. Considering this fact, our estimation system exhibits only marginal effective improvement relative to previously developed Shore hardness estimation models [14]. Further work must be done to create a truly universal compliance estimation system.
7 Conclusion
@@
We have developed a novel system for estimating the Young’s modulus of unknown contacted objects using analytical and data-driven methods. This system can create estimates online from only a single parallel grasp using camera-based tactile sensors. It has been shown to perform independent of object shape and remains robust across material types. It can precisely distinguish between compliances of soft objects, but struggles to expressively differentiate between that of harder objects.
The estimation architecture introduced in this paper could be applied for use with other camera-based tactile sensors capable of tracking depth. With more diverse tactile data, the accuracy of the system could be further improved. Given an increased ability to generalize, this compliance estimation system could be deployed alongside an intelligent high-level controller to inform decision-making in robotic manipulation scenarios by characterizing the compliance of unknown objects. This is particularly valuable when visually similar objects may have distinct material composition. As one example, it could be utilized to identify the ripeness of fruits or vegetables [1, 2, 4].
@@ Acknowledgments
The authors would like to thank Sandra Q. Liu, Jialiang (Alan) Zhao, Megha Tippur, Yuxiang Ma, and Laurence Willemet for their guidance and insight provided throughout the project’s duration.
Appendix A Method of Dimensionality Reduction (MDR)
As discussed in Section 3.2, we model contact between an unknown object and our flat, camera-based tactile sensor as a collision between an elastic sphere and an elastic half-space using Hertzian contact theory. Now, we will walk through the derivation for a constitutive relation using the method of dimensionality reduction (MDR) [58]. A mechanical diagram of contact at time $t$ with labeled variables is provided in Fig. 10.
Figure 10: Method of dimensionality reduction (MDR) diagram. An elastic sphere is colliding with an elasic half-space at time $t$. Variables of contact are labeled. The two-dimensional radial stress field $p$ is transformed into a one-dimensional force field $q_{\text{1D}}$. This transformation can be observed through comparison of top and side views.
First, we compute the maximum pressure of contact $p_{0}$ between both bodies in Eq. 6. This is a simple expression of measured force $F$ and contact radius $a$. We can substitute $a$ to create an expression dependent on the aggregate elastic compliance of the bodies $E^{*}$ [47].
$p_{0}(t)=\frac{3F(t)}{2\pi a^{2}(t)}=\frac{1}{\pi}{\left(\frac{F(t){E^{*}}^{2}}{R^{2}}\right)}^{\frac{1}{3}}$ (6)
Using this maximum pressure $p_{0}$ at timestep $t$, we create the radial pressure field of contact $p(r,t)$. This expression comes directly from Hertzian contact theory [47].
$p(r,t)=p_{0}(t)\sqrt{1-\frac{r^{2}}{a^{2}(t)}}$ (7)
Now, we will apply the method of dimensionality reduction (MDR). This technique is used to solve the stress field for surface deformation based on the assumption that contact geometries are axisymmetric. By utilizing this method, we are able to reconstruct the deformation of the sensor surface. We begin by transforming coordinates of the pressure field into one-dimensional force density $q_{\text{1D}}$ through Eq. 8.
[…]
From here, we apply a constitutive relationship on the sensor gel pad to get a one-dimensional displacement of the gel surface $u_{\text{1D}}$. This displacement is defined in the normal direction $\hat{z}$. By transforming to a single dimension, we restrict the displacement geometry to be axisymmetric about $\hat{z}$ from the center $x=0$.
$u_{\text{1D}}(x,t)=\left(\frac{1-\nu_{\text{sensor}}^{2}}{E_{\text{sensor}}}\right)q_{\text{1D}}(x,t)$ (9)
The maximum displacement will occur at $x=0$. We define this to be equivalent to our maximum measured depth $d$.
$u_{\text{1D}}(0,t)=d(t)=\left(\frac{1-\nu_{\text{sensor}}^{2}}{E_{\text{sensor}}}\right)q_{\text{1D}}(0,t)$ (10)
Then, by plugging in $x=0$, the expression for $q_{\text{1D}}$ simplifies considerably.
[…]
We solve this integral to get Eq. 13 which expresses maximum observed surface displacement $d$ in terms of normal force $F$.
[…]
$u_{\text{1D}}(0,t)=d(t)=\frac{1-\nu_{\text{sensor}}^{2}}{E_{\text{sensor}}}{\left(\frac{3{E^{*}}^{2}F(t)}{32R^{2}(t)}\right)}^{\frac{1}{3}}a(t)$ (13)
We apply this relation for every timestep $t$ over the duration of a grasp. Using all measurements of $d$, $a$, and $F$, we can apply a linear least-squares algorithm to retrieve a best fit value $\hat{E}^{*}$. Then, this estimate is used with known mechanical properties of our sensor’s silicone gel pads to compute the Young’s modulus of the unknown grasped object $\hat{E}_{\text{hertz}}$, as shown in Eq. 5.
Without the MDR, we could not have solved directly for the deformation of the gel surface, meaning we could not create an expression relating $F$ directly to $d$ in terms of Young’s modulus. Instead, we could only derive a relation for the apparent deformation of both bodies with respect to each other. This apparent deformation is equal to the displacement between both bodies, or change in gripper width $\Delta w$ in our case. This is a classic tabulated solution for Hertzian contact between a sphere and a plane [47], written as Eq. 14.
$F(t)=2E^{*}a(t)\Delta w(t)$ (14)
Notably, this model would not make use of tactile depth $d$ and relies solely on gripper width $w$ measurements. These gripper measurements exhibit a larger observed error than tactile depth measurements, which are higher-resolution and demonstrate sub-millimeter accuracy. For these reasons, we opt to use the MDR-derived solution.
Appendix B Markers
Figure 11: Markers. Tactile images from GelSight Wedge’s with and without markers. Tactile images are size 35mm by 25mm.
Markers are small black dots painted in an array on the surface of the GelSight gel pad. As the surface of the gel deforms and stretches, these markers move. Thus, we can track their relative position to map the displacement of the gel pad surface in both shear and normal directions. Markers have been shown to be effective in estimating shear and normal forces from tactile images [12, 17]. However, accurate force estimations require assumptions on contact geometry. In this work, we seek to estimate the compliance of unknown objects. Since we cannot make accurate assumptions on geometry, we measure forces through an external force gauge rather than through our GelSight sensors. In any case, we can use markers to increase the amount of dynamic information encoded in a tactile image. Without them, we could not perceive shear displacement.
In our experiment, we have created one GelSight Wedge [17] with markers painted on the surface. This was done to investigate how beneficial markers may be for estimating a dynamic property of contact like Young’s modulus. The difference between tactile images with or without markers for two objects is displayed in Fig. 11. These markers help track surface displacement, but they may occlude fine features of texture, like screw threads.
Appendix C Processing Grasp Data
We collect data over a set number of automated parallel grasps for every object in our dataset. Normal force $F$, width $w$, and maximum depth $d$ are plotted for an example grasp of a rigid sphere in Fig. 12a. Grasps are executed at nearly constant velocity to a peak force of approximately 30 N. These measurements are clipped to the loading sequence based on force thresholding. Once force passes an empirically-defined threshold of 0.75 N, we consider the object to be in-contact with the GelSight sensor and define this as timestep $t=0$. We clip data from this point until the peak recorded force is reached.
As expected in an elastic regime, we observe that force and depth measurements are high correlated from both GelSight sensors. Raw collected data normalized per grasp is plotted in Fig. 12b. Measurements are normalized according to their peak value over each respective grasp. These measurements are randomly sampled from grasps across the entire dataset. All measurements outside of 25% correlation bounds are discarded. Over 90% of measured points lie within these bounds for both sensors. We observe that there is no distinguishable difference between measurements collected from one sensor to the other.
Figure 12: Grasping data. (a) Grasp data is plotted for an example grasp of a rigid sphere. The chosen loading sequence from initial contact to peak force is highlighted in gray. (b) Normalized force and depth measurements from all collected grasps are plotted. The correlation coefficient between these measurements is computed as 0.84 and bounds are placed at 25% error mark.
Appendix D Comparison Against Shore 00 Hardness Baseline
Previous works have developed learning-based methods to estimate Shore hardness using camera-based tactile sensors without the incorporation of analytical modelling [40, 14, 5, 15, 1]. Notably, Yuan et al. introduced an architecture to estimate Shore 00 hardness of various sample geometries using tactile images captured from a GelSight sensor [14]. To demonstrate the performance of our novel hybrid estimation system, we directly compare it against results from Yuan et al.
As discussed in Section 6, we can constrain our dataset to predict compliance for only rubber objects. There are 82 rubber objects in our dataset of 285 total objects. The compliance of these rubber objects aligns with the Shore 00 hardness scale, which spans approximately from a Young’s modulus of 6 kPa to 2 MPa. We can directly evaluate and compare predictions of Young’s modulus with Shore 00 hardness by utilizing established models that convert between the two measures [53].
Using the rubber subset of our dataset, we replicate experiments conducted by Yuan et al. and compare their quoted results against our own [14]. We use RGB tactile images with markers and analytical estimates as the input to our model. First, we train an instance of our hybrid estimation system to predict compliance using only simple shapes such as spheres, cylinders, and rectangular objects. This model is validated across seen shapes with unseen compliances. Predictions from our model are subsequently converted to Shore 00 hardness to calculate error over the validation set. Second, we train another instance of our hybrid model and validate it across unseen objects with arbitrary or irregular shapes. It is important to note that Yuan et al. evaluate over hand-collected data for the latter experiment, whereas all of our data collection is performed on robot hardware. Results from our experiments are presented in Table 2, where root mean-squared error (RMSE) is reported in units of Shore 00 hardness.
Method Dataset $R^{2}$ RMSE
Shore 00 Hardness Simple Shapes 0.85 10.3
Model [14] Arbitrary Shapes 0.39 18.2
Top-1 Hybrid NN Simple Shapes 0.69 5.9
(Our Method) Arbitrary Shapes 0.37 14.3
Table 2: Shore 00 Hardness Prediction Error Comparison
Our estimation system exhibits decreased error in both experiments across both simple and complex shapes, though the margin of improvement is modest. These results suggest that our novel hybrid estimation architecture offers some performance enhancement over previously developed learning-based approaches for predicting the compliance of soft rubbers. We can likely attribute this decrease in error to the incorporation of analytical modeling.
Appendix E Dataset Diversity & Performance
The diversity of the collected dataset is characterized through categories of material and shape in Tables 4 and 4. As intended, the most common material of objects in the dataset is rubber, which includes 28.6% of the objects in the dataset. This is done to bias the dataset toward the Young’s modulus of our sensor’s gel pads. There are significant percentages of metal and plastic objects, making up 25.1% and 24.0% of the dataset respectively. Over 50% of objects in the dataset have irregular shapes. This is done to diversify training data in an attempt and learn Young’s modulus independent of object shape or contact profile. Still, there are substantial percentages of simple shapes for performance evaluation.
Table 3: Dataset Material Distributions
Material Percent (%)
Rubber 28.6%
Metal 25.1%
Plastic 24.0%
Wood 8.4%
Foam 6.3%
Other 7.6%
Table 4: Dataset Shape Distributions
Shape Percent (%)
Cylinder 20.2%
Sphere 11.5%
Rectangular 15.3%
Hexagonal 2.1%
Irregular 50.9%
Results of investigated methods on unseen objects across shape and material categories are compared in Fig. 13. We observe that neural network methods give the lowest log10 error and accuracy in nearly every category. Moreover, the accuracy increases and error decreases with the complexity of analytical models. Our Hertzian method performs best over spheres, which is expected given that this model assumes spherical contact geometry. Most critically, it is shown that our trained model can learn Young’s modulus independent of shape, given that prediction accuracy and error remain consistent across different object shapes. The model performs robustly across different categories of materials as well.
Figure 13: Estimation methods performance. Log10 error and accuracy are plotted for each estimation method across dataset categories of shape and material. Neural network results are from top-10 trained models evaluated on unseen objects.

### v3

@@ IV Physical Dataset
A novel physical dataset of 285 common household objects was gathered to train and validate methods. For comparison, this is nearly 4 times larger than the YCB dataset, which includes only 77 objects [40]. Additionally, each object was required to have approximately uniform material composition to ensure that an accurate Young’s modulus could be defined for the object and used to label data for supervised training. Furthermore, chosen objects for training must not exhibit viscoelastic behavior. A subset of total dataset objects is displayed in Fig. 4. Some objects, such as bananas and other real fruit, are not used for model training.
@@
Figure 4: Physical dataset. (a) Selected objects from our physical dataset are displayed. (b) Custom objects are molded out of silicone to replicate the shape of non-rubber objects in the dataset.
@@
To evaluate estimation algorithms’ ability to generalize, diversity of shape and material were prioritized in acquiring objects. We mold 20 custom objects out of silicone to replicate the exact shape of objects in the dataset that are made of much harder materials. Molded objects are displayed in Fig. 4b. Training on data collected across object replicas of different materials may help prevent overfitting to shapes and textures. The dataset’s composition is detailed in Tables II and II.
TABLE I: Dataset Materials
Material Percent (%)
Rubber 28.6%
Metal 25.1%
Plastic 24.0%
Wood 8.4%
Foam 6.3%
Other 7.6%
TABLE II: Dataset Shapes
Shape Percent (%)
Cylinder 20.2%
Sphere 11.5%
Rectangular 15.3%
Hexagonal 2.1%
Irregular 50.9%
As objects were gathered, they were labeled with their respective Young’s moduli. Hard objects, made of plastics, wood, glass, and metals, were labeled by referencing an engineering database of known material properties [41]. This was possible given that all selected hard objects are made of known materials. While variations due to manufacturing conditions or other factors are possible, these labels are expected to be relatively accurate with respect to our vast material domain.
Soft objects, made of rubber and foam, were labeled by converting Shore hardness measurements to Young’s modulus using Gent’s model and other conversion methods [28, 42, 43]. Each soft object was indented with a durometer 10 times and the median measurement was taken. Conversion models provide a reliable measure of object compliance on a logarithmic scale, making them a widely used approximation in material characterization [3]. To further validate assigned mechanical properties, 7 rubber and foam objects, including our sensor’s gel pad, were probed with a Dynamic Mechanical Analyzer (TA Instrument Q850) over a force range of 0 to 10 N at a rate of 2 N/min. Results from this testing are plotted in Fig. 5 confirming the accuracy of our labeling techniques. In total, the dataset includes materials ranging from foam to steel, with Young’s moduli from 5.0 kPa to 250 GPa. For reference, our sensor’s Young’s modulus is 0.275 MPa.
Figure 5: Dynamic Mechanical Analyzer results. Compressive stress-strain curves recorded on samples of selected foam and rubber objects from our dataset. Dotted lines respresent expected linear curves from labeled Young’s modulus $E$ for each object, where moduli were approximated via Shore hardness.
@@
Figure 6: Hardware for data collection and example tactile data. (a) GelSight Wedge sensors are attached to Franka Panda gripper with embedded force gauge. (b) Tactile images and depth reconstruction from a GelSight Wedge sensor without markers over a set of example objects. Shown frames are all from peak pressing force of 30 N. All displayed tactile images are size 35mm by 25mm.
V Data Collection & Training
V-A Hardware Setup
GelSight Wedge sensors [16] were rigidly attached to a Franka Panda parallel robotic gripper with 3D-printed fingers, as depicted in Fig. 6a. Fingers were printed using FormLabs Tough 2000 resin. Although we used GelSight sensors for our experiment, other camera-based tactile sensors capable of tracking contact depth in high-resolution would be suitable. One sensor was inscribed with black dot markers to help enhance the dynamic information in a tactile image by tracking surface displacement [8, 16]. Data from only one sensor is required for estimation, but we collected data from both sensors simultaneously to later assess the necessity of markers in estimating Young’s modulus. Additionally, we assume symmetry between fingers in our analytical models.
Normal force $F$ was measured through an HX711 force gauge in the left finger. Gripper position was obtained directly from control software. To improve accuracy, measured force was used to correct gripper position readings by compensating for finger deflection with a fitted linear model.
V-B Data Collection & Training
Data Collection: Automated parallel grasps were executed on objects in fixed position. For each object in our dataset, a set number of grasps were recorded. Objects were re-oriented between each grasp to diversify potential antipodal grasp locations. During grasping, tactile images from GelSight sensors, normal forces, and gripper widths, were recorded at 30 Hz. Measurements were synchronized by shifting for latency between cameras. Grasping was executed at constant velocity of 3.75 cm/s and held at closed position for 0.5 s. Peak normal force varied across each grasp, up to 30 N. Raw data collected from grasps of spherical objects is plotted in Fig. 3b.
Tactile and depth images at grasping force of 30 N for a set of example objects are depicted in Fig. 6b. As expected from Eq. 6, peak depth is higher for objects of the same shape with higher modulus. Additionally, textures are more finely captured for rigid objects with higher modulus.
Data Preprocessing: Data recorded from grasping objects was clipped to include only the loading sequence, spanning from initial contact to peak force. Misaligned grasps where loading is non-monotonic were disregarded. A fixed number of frames ($N_{\text{frames}}=3$) were sampled equidistantly in time across the loading sequence. To enhance training diversity, multiple sets of these samples were created from each grasp by shifting the starting point by one timestep. As illustrated in Fig. 2, the sampled tactile images, along with analytical estimates, normal force $F$, and width $w$ measurements, were sent as inputs to our multi-tower architecture. Additionally, the images undergo random horizontal and vertical flipping for data augmentation.
@@ V-B Data Collection & Training
Training: We trained our model for 80 episodes over 4,000 input grasps. In total, these grasps include over 12,000 tactile images. Training was performed end-to-end without any pretraining. Output labels were normalized on a log10 scale given the range of Young’s moduli in our dataset. Final sigmoid activation function was used with an MSE loss function.
VI Results
@@
Initially, we train an instance of our model using only rubber objects from our dataset. From Eq. 7, it is expected to be easier to discern fine differences in compliance between rubber objects, as their Young’s modulus is closest to that of our sensor. Estimations across unseen objects for this model are plotted in Fig. 7. The model demonstrates 100% prediction accuracy within an order of magnitude of ground truth Young’s modulus, with an average log10 error of 0.36.
Figure 7: Young’s modulus predictions for rubber objects only. Results for model trained on only rubber objects over 80 randomly sampled grasps of unseen objects. Predictions in the gray region are considered sufficiently accurate within an order of magnitude of ground truth.
@@
Figure 8: Young’s modulus predictions and model performance. (a) Results for learned and analytical estimation methods are plotted, using predictions from 80 randomly sampled grasps per material category. Predictions in the gray region are considered sufficiently accurate. (b) Log10 error and accuracy are plotted across dataset categories of shape and material. Neural network results are from top-10 trained models evaluated on unseen objects.
@@
TABLE III: Estimation Architecture Ablation Study
Furthermore, we utilize this reduced-domain training configuration to evaluate our architecture against previously-developed Shore hardness estimation models. We can only compare between predictions in this domain because previously-developed models have only been trained for use with rubber objects. Converting our model’s predictions from Young’s modulus to Shore 00 hardness [43], we find that our model reduces root mean square error (RMSE) compared to Yuan et al. [12] from 10.3 to 5.9 over simple-shaped objects and from 18.2 to 14.3 over arbitrarily-shaped objects, in units of Shore hardness. This observed improvement is likely due to our method’s integration of analytical estimates alongside tactile image-based CNN models, whereas previously methods were fully data-driven.
Subsequently, we trained our neural network over the entire collected dataset and compare against analytical baselines. Predictions across unseen objects for each method are plotted in Fig. 8a. A traditional Hooke’s Law method without tactile sensing fails to distinguish between material types. Simple elastic and Hertzian models exhibit stronger performance, but remain less than 30% accurate. Our hybrid learned model showcases dramatic improvements over all methods. Errors across all models may be partially attributed to potential inaccuracies in our dataset labelling process.
In Table III, performance is compared between different ablations of our estimation methods and inputs. Performance metrics were computed for grasps of both seen and unseen objects. We evaluated models on a logarithmic scale given the range of Young’s moduli in our dataset. We computed error as the average difference between prediction and ground truth on a log10 scale. Accuracy was computed based on the percentage of predictions within the same order of magnitude of ground truth labels.
From the ablation study, we observe that our hybrid model out-performs purely analytical and learning-based methods. RGB tactile images with markers are shown to be most favorable for this estimation task. There is only minor differences in results between models which consider normal force $F$ and width $w$. This may be due to the fact that all objects were grasped with identical peak force. Moreover, force is implicitly apparent through tactile difference images [8].
Importantly, consideration of analytical estimates $\hat{E}_{\text{elastic}},\hat{E}_{\text{hertz}}$ is shown to improve prediction accuracy by nearly 10%. This indicates our hybrid approach to compliance estimation is favorable over purely learning-based methods, which have been previously explored [12, 14, 15, 29, 30, 31]. As shown in Fig. 8b, our model performs independent of shape and can generalize well across materials.
@@
Limitations: Our estimation system struggled to express precise differences in Young’s modulus between harder objects that stray further from the modulus of our sensor (i.e. plastic, wood, and metal objects). These results are expected given that harder objects are rigid relative to our soft sensor. Moreover, the resolution of our tactile sensor is not high enough to detect minute differences in deformation between this range of compliances. Future experiments could be performed with harder sensor gels to verify if this corresponds in an increased ability to sense the compliance of harder objects. Additionally, alternative analytical models that preport to better model the non-linear, volume-conserving behavior of silicone could be explored. However, this limitation may represent an intrinsic bottleneck in the precision of depth-tracking in camera-based tactile sensors.
VII Conclusion
@@
We have developed a novel system for estimating the Young’s modulus of unknown contacted objects using analytical and data-driven methods. This system can create online estimates from only a single parallel grasp using camera-based tactile sensors. It has been shown to perform independent of object shape and remains robust across material types. It can precisely distinguish between compliances of soft objects and is more accurate over this material domain than the current state-of-the-art [12]. Moreover, by parametrizing compliance via Young’s modulus, we are able to produce estimates on an absolute scale over many material categories. This generality is demonstrated through our diverse physical dataset.
Given an ability to reliably produce estimations of compliance on a universal scale across different material regimes, this estimation system could be used alongside haptic displays to enhance teleoperation interfaces [3]. With better tactile feedback to users, more performative policies for general robotic tasks and remote surgical operation could be developed. The estimation system has already been successfully deployed to measure compliance in real-time [44].
@@ ACKNOWLEDGMENTS
The authors would like to thank Megha Tippur, Sandra Q. Liu, and Yuxiang Ma for their valuable insights and guidance.